The base of a rectangle is (x+2) cm and the height is (x-3) cm. If the area is 24 cm2, find the base and height.
Answer by amoresroy(361)

Area of rectangle = base x height
Where x+2 = the base
x-3 = the height

(x+2)(x-3) = 24
x^2-x-6-24 = 0
x^2-x-30 = 0
(x-6)(x+5) = 0
When x-6 = 0
x = 6
x+2 = 6+2 = 8
x-3 = 6-3 = 3
The base is 8 cm and the height is 3 cm.